OpenStudy (anonymous):

Use 3.14 to estimate the area of the circle the circle is 16.8m 51.12 m2 221.56 m2 448.25 m2 886.23 m2

OpenStudy (nurali):

Area=pir^2 Area=3.14*16.8^2 Area=886.23 m^2

OpenStudy (nurali):

Welcome to Openstudy.